INAF queue observing, 2015 Dec 1-2

Observer: F. Cusano, A. Rossi
Telescope Operator: Geno Bechetti
Support Astronomer: Barry Rothberg (on call)
Instrument: LUCI/LBC

Summary

At the sunset the sky is clear. We lost one hour for problem in the alignment of the LUCI mask.
Problem with the TCS, we did not observe for two and half hour.

Night duration (12\xB0 el twilight) 18:20-6:00 ( 02:30-11:50 UT) = 11 h 40m

Observing time: 5 hrs 40 m (57%)

Weather time loss: (0%)

technical time loss 5 hrs (43%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI Dark     Gilli, Iovino, Rossi, Ragazzoni
Clementini PegasusIII, Pisces II B, V 45m program completed
Iovino M004XXLrev grating 200H+K+zJfilter 2h  
Guaita M0717 H2, Br_gam 0.5h, 0.5h  
Standard Rubin149 H2, Br_gam 10m  

Details

00:13 start darks
  • for Gilli and Iovino luci2.20151202.0001-0007 (300s, mer, normal, N1.8)
  • for Rossi luci2.20151202.0008-0019 (30s, ndit=2, normal, N3.75)
  • for Ragazzoni luci2.20151202.0020-0025 (300s, mer, normal, N3.75)

1:12 Switch to LBC. Slewing to Standard photometric SA92
  • dofpia at 1:14. OPS! no stars: M3 still on
  • OB for standard not executed

1:34 restart now, slewing to Clementini field
  • running copointing
  • dofpia 1:36
  • copointing 1:47
  • seeing 1"

1:51 running Pegasus_21.obs, Pisces_21.ob, Pegasus_22.ob
  • seeing 1" blue , seeing 1.2" red at 2:00, seeing 0.9" at 2:40
  • images elongated in the red, we run dofpia again at 2:06
  • start again at 2:23 Pegasus_22.ob, Pisces_22.ob, Pegasus_23.ob, Pisces_23.ob
  • images elongated at 2:35

2:42 slew to standard SA92
  • filter LBC-B: B, r, filter LBC-R: V, I
  • seeing 0.8"

2:52 switch to LUCI2

3:21 slew to Telluric for Iovino
  • preset failed : no star to guide because calibration unit in
  • sent again (3:32)
  • acquisition luci2.20151202.0026-0028
  • spectra luci2.20151202.0029-0036
  • seeing 1"

3:50 slew to the science field for Iovino
  • acquisition luci2.20151202.0037-0041 not alligned
  • seeing 0.9"
  • start again acquisition luci2.20151202.0042-46 not alligned
  • 4:32 start again acquisition luci2.20151202.0047-51 not alligned
  • 4:52 fantastic intervention by Cusano : thanks to his genius he found the right ra/dec correction
    and the stars are now alligned. luci2.20151202.0052-53
  • 4:54 Start spectra luci2.20151202.0054-0077
  • 5:42 seeing 1.1"
  • 6:26 seeing 1.0"
  • 6:48 problem with the GUI

7:22 moving to Guaita imaging H2, luci2.20151202.0078-00107
  • at the start seeing 0.9"
  • 07:30 seeing 0.6" on image
  • 07:40 seeing 0.5" !!
  • 08:08 seeing 0.4" on image

8:15 slewing to standard Rubin149, filter H2 luci2.2015120108-117

8:25 standard Rubin149, filter Br_gam luci2.2015120118-127

8:35-9:00 slewing to Guaita imaging Br_gam, problem with the telescope preset, the TO is checking in the chamber
  • 9:00 Sent again OB but LUCI not responding, problem with TCS.
  • 9:00 - 9:30 LBT staff working on the TCS problem
  • 9:32 TCS ok, preset complete, guiding not working, needs new pointing and collimation

10:02 the preset is fine, Guaita Br_gam luci2.20151202.128-157
  • seeing 0.6"
  • seeing 0.7" at 10:28

10:55 slew to telluric of Gilli program,
  • preset failed can not find the guide star 10:58
  • seeing 0.5"
  • 11:05 starting again acquisition luci2.20151202.159-161
  • spectra 162-164 satured
  • spectra 165-166 are ok!

11:27 slewing to science of Gilli; we are try to authorise LBC-B but is not working,
  • problem authorizing LUCI2+LBCb
  • same problem with LBCb+LBCr
  • try to switch to LUCI only but does not work
  • still problem with the TCS
Topic revision: r28 - 20 Dec 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback